A manufacturer estimates that its variable cost for manufacturing a given product
is given by the following expression:
C(q) = 25q2 + 2000q [$] where C is the total cost and q is the quantity produced
a. Derive an expression for the marginal cost of production
b. Derive expressions for the revenue and the profit when the widgets are sold
at marginal cost.
Answer #1

Answer

a)

MC is the change in total cost and a cange in function found by differentiation

MC=dC/dq=50q+2000

b)

the price is equal to MC

P=MC=50q+2000

TR=P*Q=50q^2+2000q

profit =TR-TC=50q^2+2000q-(25q^2+2000q)

=25q^2

the profit =25q^2